OpenStudy (anonymous):

Given that √(5-2(√6))=√a-√b, find the values of 'a' and 'b'

OpenStudy (anonymous):

\[\sqrt{5-2\sqrt{6}}=\sqrt{a}-\sqrt{b}\]

OpenStudy (anonymous):

try squaring both sides

OpenStudy (anonymous):

I am not quite sure if I am doing this right \[5-2\sqrt{6}=(\sqrt{a}-\sqrt{b})^{2}\] \[(\sqrt{a}-\sqrt{b})^{2}=(\sqrt{a}-\sqrt{b})(\sqrt{a}-\sqrt{b})\] \[5-2\sqrt{6}=a-2\sqrt{ab}+b\]

OpenStudy (anonymous):

yup looks good so far

OpenStudy (anonymous):

oh so a+b=5 and a*b=6, solve for a and b

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

the a will be 3 and b will be 2

OpenStudy (anonymous):

except only issue would be determining which one is a and which one is b

OpenStudy (anonymous):

oh, ya

OpenStudy (anonymous):

oh wait a= higher number cant get a negative from a square root

OpenStudy (anonymous):

ok, thank you so much!!!!